OpenStudy (anonymous):

Derivative of: f(x) = (e^7x)(sin(8x)) g(x) = (x^6)(sin(9x)) h(t) = 8(tan(t^6))

OpenStudy (amistre64):

product rule and chain rule will get you thru it

OpenStudy (amistre64):

if there is a specific question you have about it, just ask, otherwise this just seems to come across as you wanting a free answering service

OpenStudy (anonymous):

f(x) = (e^7x)(sin(8x)) f'(x) = (e^7x)(sin8x)' + (sin8x)(e^7x)' for (d/dx)(sinnx)=dsinnx/dnx = ncosnx for d/dx (e^nx) = de^nx/dnx = ne^nx you can then find the ans Question (2) is similar to quesiton (1) h(t) = 8(tan(t^6)) dtanx/dx=sec^x dt^ndt = nt^(n-1) then, you can get the ans Try it yourself

OpenStudy (anonymous):

The answers I got were: F'(x) = 56(e^(7x))(cos(8x)); g(x) = 54(x^5)(cos(9x)); and h(t) = 6(t^5)(sec(t^6))^2... and apparently they're incorrect. I'm well aware amistre that I have to do product and chain rule and have attempted that... you don't have to be obnoxious to people looking for a little help.

OpenStudy (anonymous):

amendment : dtanx/dx=sec^2 x

OpenStudy (amistre64):

actually, i do in order to evaluate your responses and intention in order to police the site appropriately.

OpenStudy (anonymous):

f'(x) = (e^7x)(sin8x)' + (sin8x)(e^7x)' = 8e^7x cos 8x <--- dsinx=cosx (don't forget to d8x which is inside sin) + 7sin8x e^7x <--- de^7x =7e^7x (don't forget d7x=7) =e^7x( 8cos8x+7sin8x) <--- take out the common factor

OpenStudy (anonymous):

g(x) = (x^6)(sin(9x)) g'(x) = (x^6)'(sin(9x))+(x^6)(sin(9x))' =6x^5 sin9x +9 x^6 cos9x = 3x^5 ( 2sin9x+3xcos9x) This is similar to f'(x)

OpenStudy (anonymous):

oh okay.. tnx CoCoTsoi.. I see what I missed.. thanks again

OpenStudy (anonymous):

I got it from here

OpenStudy (anonymous):

h(t) = 8(tan(t^6)) h'(t) = 8 sec^2 (t^6) 6t^5 = 48t^5 sec^2 (t^6)

OpenStudy (amistre64):

to mirror coco [fg]' = f'g+fg' f(x) = (e^7x)(sin(8x)) f'(x) = (e^7x)'(sin(8x)) + (e^7x)(sin(8x))' f'(x) = 7(e^7x)(sin(8x)) + 8(e^7x)(cos(8x)) f'(x) = e^7x(7sin(8x) + 8cos(8x))

OpenStudy (anonymous):

I am glad to help you :D